OpenStudy (angelwings996):

What is the solution of the equation? (x - 2)^3/4 = 8

OpenStudy (abb0t):

multiply by 4 and cube everything. then add 2 to solve for x.

OpenStudy (angelwings996):

\[(x - 2)^{3/4} = 8 \] is the equation

OpenStudy (abb0t):

That's the same thing as: \[\sqrt[4]{a^3}\]

OpenStudy (angelwings996):

@abb0t will there only be one solution?

OpenStudy (anonymous):

((x - 2)^1/4)^3 = 8 =2^3 it powers are same then you can compare the bases i.e (x - 2)^1/4 = 2 then multiply the both sides with ^4 then you will get x-2=16 x=18

OpenStudy (anonymous):

This one won't have multiple solutions.

OpenStudy (anonymous):

its like (x-2)^3 = 8^4 giving you x = 18. And only 1 root

OpenStudy (angelwings996):

@mominanisar Thanks!

OpenStudy (angelwings996):

Could any of you help me with one more please?

OpenStudy (anonymous):

no problem

OpenStudy (angelwings996):

You live near a bridge that goes over a river. The underside of the bridge is an arch that can be modeled with the function y = -0.000471x^2 + .837x where x and y are measured in feet. How high above the river is the bridge (the top of the arch)? Hw long is the section of bridge above the arch?

OpenStudy (angelwings996):

@mominanisar @Saikam could either of you help me please?

OpenStudy (angelwings996):

It is multiple choice so I can list them

OpenStudy (anonymous):

yes i will if i am able to

OpenStudy (angelwings996):

Okay, ill give you the choices

OpenStudy (anonymous):

actaully, the length of the section of bridge and its height above the river are dependant on each other. one cannot be determined without the other

OpenStudy (angelwings996):

a. The bridge is about 1,777.07 ft. above the river, and the length of the bridge above the arch is about 888.54 ft. b. The bridge is about 371.85 ft. above the river, and the length of the bridge above the arch is about 1,777.07 ft. c. The bridge is about 371.85 ft. above the river, and the length of the bridge above the arch is about 888.54 ft. d. The bridge is about 1,777.07 ft. above the river, and the length of the bridge above the arch is about 371.85 ft.

OpenStudy (angelwings996):

@Saikam That's what everyone is saying, so Im not able to find the answer anywhere. Im not even sure how you would figure this out

OpenStudy (anonymous):

i agree with @Saikam

OpenStudy (angelwings996):

Then would I have to take an educated guess on this?

OpenStudy (anonymous):

no. we must consider each option and verify. take the length of bridge given in the first option, then check if it satisfies the given height. Then take the values from second option, and so on... one (or more) of them might satisfy. can u try? or u want me to help?

OpenStudy (angelwings996):

Could you help me please?

jimthompson5910 (jim_thompson5910):